# Similarity of Neural Architectures using Adversarial Attack Transferability

### v3

@@ 5.2 Model diversity and knowledge distillation
Knowledge distillation [45] is a training method for transferring rich knowledge of a well-trained teacher network. Intuitively, distillation performance affects a lot by choice of the teacher network; however, the relationship between similarity and distillation performance has not yet been explored enough, especially for ViT. This subsection investigates how the similarity between teacher and student networks contributes to the distillation performance. There are several studies showing two contradictory conclusions; Jin et al. [51] and Mirzadeh et al. [70] showed that a similar teacher leads to better distillation performance; Touvron et al. [96] reports that distillation from a substantially different teacher is beneficial for ViT training.
@@
6 Limitations and discussions
We include three discussions related to SAT. Due to the space limit, we leave more details in Appendix. Our discussions include: (1) How adversarially trained models affect SAT score. (2) An efficient approximation of our analysis for a novel model (i.e., for a novel 70-th model for our analysis). Instead of generating adversarial attacks for all 70 models, just using attacked images from the novel model could be enough. (3) Other possible applications of SAT, e.g., a novel architecture development, analyzing the hyper-parameter sensitivity, model selection for diverse models.
@@ 7 Conclusion
We have explored similarities between image classification models to investigate what makes the model similar or diverse and whether developing and using diverse models is required. For quantitative and model-agnostic similarity assessment, we have suggested a new similarity function, named SAT, based on attack transferability demonstrating differences in input gradients and decision boundaries. Using SAT, we conduct a large-scale and extensive analysis using 69 state-of-the-art ImageNet models. We have shown that macroscopic architectural properties, such as base architecture and stem architecture, have a more significant impact on similarity than microscopic operations, such as types of convolution, with numerical analysis. Furthermore, the effectiveness of the training strategy is minor compared to model architectural designs. Finally, we have first discovered the advantages of using diverse models in ensemble or distillation strategies for a practical scenario with a large-scale training dataset and a highly complex architecture.
@@
This work is done as an internship project by J Hwang under the supervision of S Chun. S Chun initialized the project idea: understanding how different architectures behave differently by using an adversarial attack. J Hwang, S Chun, and D Han jointly designed the analysis toolbox. J Hwang implemented the analysis toolbox and conducted the experiments with input from S Chun, D Han, and J Lee. J Hwang, D Han, B Heo, and S Chun contributed to interpreting and understanding various neural architectures under our toolbox. The initial version of “model card” (Tab. A.3 and A.4) was built by J Hwang, S Park, and verified by D Han and B Heo. B Heo contributed to interpreting distillation results. All ResNet and ViT models newly trained in this work were trained by S Park. J Lee supervised J Hwang and verified the main idea and experiments during the project. S Chun and J Hwang wrote the initial version of the manuscript. All authors contributed to the final manuscript.
Supplementary Material
We include additional materials in this document, such as the network details, feature importance analysis details, spectral clustering method, training settings, and setting of knowledge distillation. We also provide experimental results of spectral clustering, computing pairwise similarities, ensemble performance, and distillation performance. We discuss some limitations and discussion points of our work, such as the effect of an efficient approximation, and possible applications.
Appendix A Details of networks used for our analysis
@@
We use 69 models in our research to evaluate the similarity between models and to investigate the impact of model diversity. In the main paper, we mark the names of models based on their research paper and PyTorch Image Models library (timm; 0.6.7 version) [107]. Tab. A.1 shows the full list of the models based on their research paper and timm alias.
@@
In Tab. A.2, we show brief information about the architectural components. The full network specification is shown in Tab. A.3 and Tab. A.4. We follow the corresponding paper and timm library to list the model features.
@@
Table A.1: Lists of 69 models and their names based on their research paper and timm library.
@@
Table A.2: Overview of model elements. We categorize each architecture with 13 different architectural components. The full feature list of each architecture is in Tab. A.3 and Tab. A.4.
Table A.3: Description of features of 69 models. “s” in “Stem layer” indicates the stride of a layer in the stem, and the number before and after “s” are a kernel size and size of stride, respectively. For example, “3s2/3/3” means that the stem is composed of the first layer having 3 $\times$ 3 kernel with stride 2, the second layer having 3 $\times$ 3 kernel with stride 1, and the last layer having 3 $\times$ 3 with stride 1.
@@
Table A.4: Description of features of 69 models. “Pooling (stem)” and “Pooling (final)” denote “Pooling at the stem” and “Pooling for final feature”, respectively. “SA”, “CW”, and “DW” means “Self-attention”, “Channel-wise”, and “Depth-wise”, respectively.
@@
Appendix B Details of feature importance analysis
@@
Appendix C Spectral clustering details
Figure C.1: Spectral features of 69 architectures. The $K$-th largest eigenvectors of the Laplacian matrix of the pairwise similarity graph of 69 architectures are shown ($K=10$ in this figure). Rows and columns are sorted by the clustering index in Tab. 2. We denote the model name in timm for each row and column.
We use the normalized Laplacian matrix to compute the Laplacian matrix. We also run K-means clustering 100 times and choose the clustering result with the best final objective function to reduce the randomness by the K-means clustering algorithm.
We visualize the pairwise distances of the spectral features (i.e., $K$-largest eigenvectors of $L$) of 69 architectures with their model names in Fig. C.1. This figure is an extension of Fig. 4(a) of the main paper, now including the model names. Note that rows and columns of Fig. C.1 are sorted by the clustering results. Fig. C.1 shows block diagonal patterns, i.e., in-cluster similarities are large while between-cluster similarities are small.
Appendix D Pairwise similarities
@@
Figure D.1: Pairwise similarity among 69 models. Rows and columns are sorted by the clustering index in Tab. 2. ($n$, $n$) component of pairwise similarity is close to 4.6 ($\log 100$) because the attack success rate is almost 100% when a model used for generating adversarial perturbation and attacked model are the same.
@@
Fig. D.1 indicates the pairwise similarity among 69 models. We can observe a weak block pattern around clusters, as also revealed in Fig. C.1.
Appendix E Training setting details for Sec. Training strategy and similarity score
@@
We train 21 ResNet-50 models and 16 ViT-S from scratch individually by initializing each network with different random seeds. We further train 28 ResNet-50 models by randomly choosing learning rate ($\times$0.1, $\times$0.2, $\times$0.5, $\times$1, $\times$2, and $\times$5 where the base learning rate is 0.1), weight decay ($\times$0.1, $\times$0.2, $\times$0.5, $\times$1, $\times$2, and $\times$5 where the base weight decay is 1e-4), and learning rate scheduler (step decay or cosine decay). Similarly, we train 9 ViT-S models by randomly choosing learning rate ($\times$0.2, $\times$0.4, and $\times$1 where the base learning rate is 5e-4) and weight decay ($\times$0.2, $\times$0.4, and $\times$1 where the base weight decay is 0.05). Note that the DeiT training is unstable when we use a larger learning rate or weight decay than the base values. Finally, we collect 22 ResNet-50 models with different training regimes: 1 model with standard training by PyTorch [74]; 4 models trained by GluonCV [37]111gluon_resnet50_v1b, gluon_resnet50_v1c, gluon_resnet50_v1d, and gluon_resnet50_v1s from timm library.; a semi-supervised model and semi-weakly supervised model on billion-scale unlabeled images by Yalniz et al. [111]222ssl_resnet50 and swsl_resnet50 from timm library.; 5 models trained by different augmentation methods (Cutout [26], Mixup [117], manifold Mixup [103], CutMix [114], and feature CutMix [^3] We use the official weights provided by https://github.com/clovaai/CutMix-PyTorch.; 10 optimized ResNet models by [108] [^4] We use the official weights provided by https://github.com/rwightman/pytorch-image-models/releases/tag/v0.1-rsb-weights. We also collect 7 ViT-S models with different training regimes, including the original ViT training setup [28]555deit_small_patch16_224, vit_small_patch16_224, deit_small_distilled_patch16_224, and deit3_small_patch16_224 from timm library., a stronger data augmentation setup in the Deit paper [96]footnotemark: , the training setup with distillation [96]footnotemark: , an improved DeiT training setup [98]footnotemark: , and self-supervised training fashions by MoCo v3 [13] [^6] We train the ViT-S model by following https://github.com/facebookresearch/moco-v3, MAE [42] [^7] We train the ViT-S model by following https://github.com/facebookresearch/mae and BYOL [35] [^8] We train the ViT-S model by following https://github.com/lucidrains/byol-pytorch. We do not use adversarially-trained networks because the adversarial training usually drops the standard accuracy significantly [100].
Appendix F Ensemble performance
Fig. F.1 shows the ensemble performance by ImageNet top-1 error (lower is better). It has the same tendency as the error reduction rate reported in Fig. 6 of the main paper.
Figure F.1: Model diversity by architecture difference and ImageNet top-1 error ensemble performance. We report ImageNet top-1 error by varying the number of ensembled models ($N$) and the diversity of the ensembled models. The diversity is controlled by choosing the models from $k$ different clusters (larger $k$ denotes more diversity).
Appendix G Knowledge distillation
G.1 Training setting details for Sec. Model diversity and knowledge distillation
@@ Appendix G Knowledge distillation
G.2 Teacher accuracy and distillation performance

### v4

@@ 5.2 Model Diversity and Knowledge Distillation
Knowledge distillation (KD) [47] is a training method for transferring rich knowledge of a well-trained teacher network. Intuitively, KD performance affects a lot by choice of the teacher network; however, the relationship between similarity and KD performance has not yet been explored enough, especially for ViT. This subsection investigates how the similarity between teacher and student networks contributes to the distillation performance. There are several studies showing two contradictory conclusions; Jin et al. [54] and Mirzadeh et al. [73] showed that a similar teacher leads to better KD performance; Touvron et al. [100] reports that distillation from a substantially different teacher is beneficial for ViT.
@@
6 Discussion
In Appendix 0.G, we describe more discussions related to SAT. We first propose an efficient approximation of SAT when we have a new model; instead of generating adversarial samples from all models, only generating adversarial samples from the new model can an efficient approximation of SAT (Sec. 0.G.1). We also show that SAT and the same misclassified samples have a positive correlation in Sec. 0.G.2. Sec. 0.G.3 demonstrates that we can estimate the similarity with a not fully trained model (e.g., a model in an early stage). Finally, we describe more possible applications requiring diverse models (Sec. 0.G.4).
@@ 7 Conclusion
We have explored similarities between image classification models to investigate what makes the model similar or diverse and whether developing and using diverse models is required. For quantitative and model-agnostic similarity assessment, we have suggested a new similarity function, named SAT, based on attack transferability demonstrating differences in input gradients and decision boundaries. Using SAT, we conduct a large-scale and extensive analysis using 69 state-of-the-art ImageNet models. We have shown that macroscopic architectural properties, such as base architecture and stem architecture, have a more significant impact on similarity than microscopic operations, such as types of convolution, with numerical analysis. Finally, we have provided insight into the ML applications using multiple models based on SAT, e.g., model ensemble or knowledge distillation. Overall, we suggest using SAT to improve methods with multiple models in a practical scenario with a large-scale training dataset and a highly complex architecture.
Acknowledgement
We thank Taekyung Kim and Namuk Park for comments on the self-supervised pre-training. This work was supported by an IITP grant funded by the Korean Government (MSIT) (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and by the Yonsei Signature Research Cluster Program of 2024 (2024-22-0161).
@@ Acknowledgement
This work is done as an internship project by J Hwang under the supervision of S Chun. S Chun initialized the project idea: understanding how different architectures behave differently by using an adversarial attack. J Hwang, S Chun, and D Han jointly designed the analysis toolbox. J Hwang implemented the analysis toolbox and conducted the experiments with input from S Chun, D Han, and J Lee. J Hwang, D Han, B Heo, and S Chun contributed to interpreting and understanding various neural architectures under our toolbox. The initial version of “model card” (Tab. 0.C.3 and 0.C.4) was built by J Hwang, S Park, and verified by D Han and B Heo. B Heo contributed to interpreting distillation results. All ResNet and ViT models newly trained in this work were trained by S Park. J Lee supervised J Hwang and verified the main idea and experiments during the project. S Chun and J Hwang wrote the initial version of the manuscript. All authors contributed to the final manuscript.
Appendix 0.A Empirical Comparison of SAT and Other Methods
Table 0.A.1: Comparison of stability of measurements. We compare the stability of method by [90] and SAT. Stability is indicated by the standard deviation (std). The numbers in ($\cdot$) mean the sampling ratio to all possible combinations to compute the exact value. “cost” denotes relative costs compared to the total forward costs for the 50K ImageNet validation set: Somepalli et al. needs ($50\text{K}\choose 3$ = 2.1×$\times 10^{13}$ and SAT needs 50K. Here, we assume that forward and backward computations cost the same.
Somepalli et al. [90] SAT (ours)
# triplets std cost # images std cost
10 (4.8×$\times 10^{-13}$) 4.49 1.0 500 (0.01) 1.88 1.0
20 (9.6×$\times 10^{-13}$) 3.28 2.0 1000 (0.02) 1.05 2.0
50 (2.4×$\times 10^{-12}$) 1.63 5.0 2500 (0.05) 0.91 5.2
100 (4.8×$\times 10^{-12}$) 1.54 10.0 5000 (0.1) 0.77 10.2
0.A.1 Comparison with Somepalli et al. in the Variance of Similarity
Somepalli et al. [90] proposed a sampling-based similarity score for comparing decision boundaries of models. SAT has two advantages over Somepalli et al.: computational efficiency and the reliability of the results. First, SAT involves sampling 5,000 images and using 50-step PGD; the computation cost is [5,000 (sampled images) $\times$ 50 (PGD steps) + 5,000 (test to the other model)] $\times$ 2 (two models). Meanwhile, Somepalli et al. [90] sample 500 triplets and generate 2,500 points to construct decision boundaries. In this case, the total inference cost is [500 (sampled triplets) $\times$ 2,500 (grid points)] $\times$ 2 (two models), which is 4.9 times larger than SAT. Secondly, Somepalli et al. sampled three images of different classes. As the original paper used CIFAR-10 [59], 500 triplets can cover all possible combinations of three classes among the ten classes (($10\choose 3$ = 120 $<$ 500). However, it becomes computationally infeasible to represent all possible combinations of three classes among many classes (e.g., ImageNet needs ($1000\choose 3$ = 164,335,500, 130 times greater than its training images). Also, we find that the similarity of [90] is unreliable when the number of sample triplets is small. In Tab. 0.A.1, we calculate the similarity scores between ConvNeXt-T [69] and Swin-T [68] from ten different sets with varying sample sizes. SAT exhibits significantly better stability (i.e., low variances) than Somepalli et al. Note that we use our similarity measurement as the percentage degree without the logarithmic function and control the scale of samples to maintain similar computation complexity between SAT and the compared method for a fair comparison.
0.A.2 Ensemble Performance Comparison of SAT and Other Methods
The purpose of our study is to provide a new lens for model similarity through adversarial attack transferability. Since we do not have the ground truth of the “similarity” between architectures, comparing different similarity functions is not really meaningful. Instead, we indirectly compare SAT, Somepalli et al. [90] and naive architecture feature-based clustering on the ensemble benchmark. More specifically, the naive architecture clustering is based on our architecture features proposed in Sec. 4.1; we apply the K-means clustering algorithm to get clusters. We note that the other comparison methods cannot be applied due to the expensive computations. The results are shown in Tab. 0.A.3 and Tab. 0.A.3.
Table 0.A.2: Somepalli et al.
less diverse $\leftarrow$ # of clusters $\rightarrow$ more diverse
1 2 3 4 5
# of models 2 7.54 7.79
3 10.85 11.03 11.13
4 12.63 12.79 12.87 12.93
5 13.74 13.89 13.95 14.01 14.05
Table 0.A.3: Arc-based clustering
less diverse $\leftarrow$ # of clusters $\rightarrow$ more diverse
1 2 3 4 5 rand SAT
# of models 2 7.60 7.80 7.78 7.84
3 11.02 11.08 11.12 11.11 11.20
4 13.03 12.92 12.89 12.90 12.90 13.00
5 14.37 14.13 14.03 14.00 14.00 14.01 14.11
In the tables, SAT is the best-performing model similarity score on the ensemble task. We also tried to Tramèr et al. [103] in the same setting, but Tramèr et al. [103] often failed to converge and show very small differences. Hence, we couldn’t use Tramèr et al. [103] for measuring all 69 arches used in the paper. Instead, as we reported in the main paper, we compare Tramèr et al. [103] and other model similarity variants on 8 architectures used in Fig. 0.B.1. Among the candidate methods, we observe that SAT shows the strongest correlation with the ensemble performance using two models.
Appendix 0.B Discussions
0.B.1 Robustness of SAT to the choice of the attack methods.
(a) PGD vs. Autoattack
(b) PGD vs. MIFGSM
(c) PGD vs. VMIFGSM
(d) PGD vs. Low Freq. PGD
(e) PGD vs. Patchfool
(f) PGD vs. BIA
Figure 0.B.1: Effect of different adversarial attack methods to SAT. The trend line and its 90% confidence interval are shown. We show the relationship between our SAT using PGD [70] and SAT using other attacks, (a) Autoattack [23] (b) MIFGSM [29](c) VMIFGSM [112] (d) low-frequency PGD [38] (e) Patchfool attack [33], and (f) BIA [129].
Our goal is not to design an attack-free method but to show the potential of using adversarial attack transferability (AAT) for measuring quantitative model similarity. However, we have checked that SAT scores are robust to the choice of the attack methods, even for stronger attacks, such as AutoAttack [23], attacks designed for enhancing AAT, such as MIFGSM [29] and VMIFGSM [112], low-frequency targeted attacks, such as low-frequency PGD [38], method-specific attacks, such as PatchPool [33], or generative model-based attacks, such as BIA [129]. We sample 8 representative models among 69 models for testing the effect of attacks on SAT; ViT-S, CoaT-Lite Small, ResNet-101, LamHaloBotNet50, ReXNet-150, NFNet-L0, Swin-T and Twins-pcpvt. Fig. 1(a) shows the high correlation between SAT scores using PGD and Autoattack; it shows a correlation coefficient of 0.98 with a p-value of $1.43\times 10^{-18}$. For testing the Patchfool attack, we only generate adversarial perturbations on ViT-S and get attack transferability to all other models (68 models) because it only targets Transformers. Fig. 1(b) and 1(c) show similar results: SAT shows consistent results even for the attacks designed for better AAT. The results show that SAT score is robust to the choice of attack methods if the attack is strong enough. Also, the low-frequency attack [38] shows a similar result, i.e., the frequency-targeted attack does not affect the similarity results. In Fig. 1(e), Patchfool also shows a high correlation compared to the PGD attack (correlation coefficient 0.91 with p-value $3.62\times 10^{-27}$). We additionally provide a result for generative model-based attack, BIA [129]. As BIA needs to train a new generative model for a different architecture, we only show the pre-trained models provided by the authors. We also get a similar result with previous results for BIA. Note that the compared attack methods are not model agnostic or computationally expensive than PGD, e.g., PatchFool needs a heavy modification on the model code to extract attention layer outputs manually, and BIA needs to train a new generator for a new architecture. As SAT shows consistent rankings across the attack methods, we use PGD due to its simplicity.
0.B.2 Impact of Adversarial Training to SAT
While our main analyses are based on ImageNet-trained models, in this subsection, we use CIFAR-10-trained models for two reasons. First, it is still challenging to achieve a high-performing adversarially trained model on the ImageNet scale. On the other hand, in the CIFAR-10 training setting, a number of adversarially-trained models are available and comparable. Second, adversarial training models show lower clean accuracy than normally trained models [105]. Adversarial robustness and accuracy are in a trade-off, and there is no ImageNet model with accuracy aligned with our target models yet.
We choose five adversarial training ResNet-18 from the AutoAttack repository [23] and measure SAT using the models. The average SAT between adversarial training models is 3.15, slightly lower than the similarity score with different training strategies for ImageNet ResNet-50 (3.27 – See Tab. 2). In other words, we can confirm that different adversarial training methods make as a difference as different training techniques.
Appendix 0.C Details of Architectures Used in the Analyses
@@
We use 69 models in our research to evaluate the similarity between models and to investigate the impact of model diversity. In the main paper, we mark the names of models based on their research paper and PyTorch Image Models library (timm; 0.6.7 version) [114]. Tab. 0.C.1 shows the full list of the models based on their research paper and timm alias.
@@
We show brief information of the architectural components in Tab. 0.C.2. The full network specification is shown in Tab. 0.C.3 and Tab. 0.C.4. We follow the corresponding paper and timm library to list the model features.
Table 0.C.1: Lists of 69 models and their names based on their research paper and timm library.
@@
Table 0.C.2: Overview of model elements. We categorize each architecture with 13 different architectural components. The full feature list is in the Appendix.
@@
Table 0.C.3: Description of features of 69 models. “s" in “Stem layer" indicates the stride of a layer in the stem, and the number before and after “s" are a kernel size and size of stride, respectively. For example, “3s2/3/3" means that the stem is composed of the first layer having 3 $\times$ 3 kernel with stride 2, the second layer having 3 $\times$ 3 kernel with stride 1, and the last layer having 3 $\times$ 3 with stride 1.
@@
Table 0.C.4: Description of features of 69 models. “Pooling (stem)" and “Pooling (final)" denote “Pooling at the stem" and “Pooling for final feature", respectively. “SA", “CW", and “DW" means “Self-attention", “Channel-wise", and “Depth-wise", respectively.
@@
Appendix 0.D Feature Importance and Clustering Details
0.D.1 Feature Importance Analysis Details
@@
Figure 0.D.1: Pairwise similarity among 69 models. Rows and columns are sorted by the clustering index in Tab. 2. ($n$, $n$) component of pairwise similarity is close to 4.6 ($\log 100$) because the attack success rate is almost 100% when a model used for generating adversarial perturbation and attacked model are the same.
0.D.2 Pairwise Similarities
@@
Fig. 0.D.1 indicates the pairwise similarity among 69 models. We can observe a weak block pattern around clusters, as also revealed in Fig. 0.D.2.
0.D.3 Spectral Clustering Details
We use the normalized Laplacian matrix to compute the Laplacian matrix. We also run K-means clustering 100 times and choose the clustering result with the best final objective function to reduce the randomness by the K-means clustering.
Figure 0.D.2: Spectral features of 69 architectures. The $K$-th largest eigenvectors of the Laplacian matrix of the pairwise similarity graph of 69 architectures are shown ($K=10$ in this figure). Rows and columns are sorted by the clustering index in Tab. 2. We denote the model name in timm for each row and column.
We visualize the pairwise distances of the spectral features (i.e., $K$-largest eigenvectors of $L$) of 69 architectures with their model names in Fig. 0.D.2. This figure is an extension of Fig. 4, now including the model names. Note that rows and columns of Fig. 0.D.2 are sorted by the clustering results. Fig. 0.D.2 shows block diagonal patterns, i.e., in-cluster similarities are large while between-cluster similarities are small.
Appendix 0.E Training Settings for Models Used in Analyses
Models with various training methods for Sec. 4.1.
@@ Appendix 0.E Training Settings for Models Used in Analyses
We train 21 ResNet-50 models and 16 ViT-S from scratch individually by initializing each network with different random seeds. We further train 28 ResNet-50 models by randomly choosing learning rate ($\times$0.1, $\times$0.2, $\times$0.5, $\times$1, $\times$2, and $\times$5 where the base learning rate is 0.1), weight decay ($\times$0.1, $\times$0.2, $\times$0.5, $\times$1, $\times$2, and $\times$5 where the base weight decay is 1e-4), and learning rate scheduler (step decay or cosine decay). Similarly, we train 9 ViT-S models by randomly choosing learning rate ($\times$0.2, $\times$0.4, and $\times$1 where the base learning rate is 5e-4) and weight decay ($\times$0.2, $\times$0.4, and $\times$1 where the base weight decay is 0.05). Note that the DeiT training is unstable when we use a larger learning rate or weight decay than the base values. Finally, we collect 22 ResNet-50 models with different training regimes: 1 model with standard training by PyTorch [79]; 4 models trained by GluonCV [39]111gluon_resnet50_v1b, gluon_resnet50_v1c, gluon_resnet50_v1d, and gluon_resnet50_v1s from timm library.; a semi-supervised model and semi-weakly supervised model on billion-scale unlabeled images by Yalniz et al. [118]222ssl_resnet50 and swsl_resnet50 from timm library.; 5 models trained by different augmentation methods (Cutout [27], Mixup [125], manifold Mixup [108], CutMix [122], and feature CutMix [^3] We use the official weights provided by https://github.com/clovaai/CutMix-PyTorch.; 10 optimized ResNet models by [115]444 We use the official weights provided by https://github.com/rwightman/pytorch-image-models/releases/tag/v0.1-rsb-weights. We also collect 7 ViT-S models with different training regimes, including the original ViT training setup [30]555deit_small_patch16_224, vit_small_patch16_224, deit_small_distilled_patch16_224, and deit3_small_patch16_224 from timm library., a stronger data augmentation setup in the Deit paper [100]footnotemark: , the training setup with distillation [100]footnotemark: , an improved DeiT training setup [102]footnotemark: , and self-supervised training fashions by MoCo v3 [15] [^6] We train the ViT-S model by following https://github.com/facebookresearch/moco-v3, MAE [44] [^7] We train the ViT-S model by following https://github.com/facebookresearch/mae and BYOL [37] [^8] We train the ViT-S model by following https://github.com/lucidrains/byol-pytorch. We do not use adversarially-trained networks because the adversarial training usually drops the standard accuracy significantly [105].
Distillation models for Sec. 5.
@@
Appendix 0.F Knowledge Distillation
0.F.1 Teacher Accuracy and Distillation Performance